OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 8 minus 2 times x all over 3 times x minus 12.? [Hint: (a−x) = − (x − a)] 2 over 3. 3 over 2. x − 4 x + 4

OpenStudy (ash2326):

\[\frac{8-2x}{3x-12}\] Did you try @jaderbrown

OpenStudy (anonymous):

yeah it's either c or d im sure

OpenStudy (ash2326):

Can you take any common term out of the numerator? \[(8-2x)\]

OpenStudy (anonymous):

2

OpenStudy (ash2326):

good, what would you get? could you bring 2 out? and write what would you get

OpenStudy (anonymous):

I'm confused, what do you mean?

OpenStudy (ash2326):

Can you bring out 2 from numerator?

OpenStudy (anonymous):

yes?

OpenStudy (ash2326):

And what would you get then? I mean the numerator

OpenStudy (anonymous):

2 over 3

OpenStudy (ash2326):

I mean like this \[8-2x\] =>\[2(4-x)\]

OpenStudy (anonymous):

alright then what?

OpenStudy (ash2326):

Now, we have to do the same thing with the denominator. Take out common factor from denominator and write the expression you get

OpenStudy (anonymous):

3(3-x)?

OpenStudy (ash2326):

you have 3x-12 Take out 3 from this. I'll simplify it a bit for you \[3\times x -3\times 4\]

OpenStudy (anonymous):

alright now what?

OpenStudy (ash2326):

Please bring 3 out from this \[3x-3\times 4\]

OpenStudy (anonymous):

3x-12

OpenStudy (ash2326):

@jaderbrown We have to bring a common 3 out from both ther terms \[3x-12\] \[(3\times x -3\times 4)\] \[3(x-4)\] Do you understand this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now what

OpenStudy (ash2326):

Now we have \[\frac{2(4-x)}{3(x-4)}\] do you see any common factor?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

x+4

OpenStudy (ash2326):

We can write \[(4-x)=-(x-4)\] so \[\frac{2(4-x)}{3(x-4)}=\frac{-2(x-4)}{3(x-4)}\] do you follow this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

So we see \[\frac{-2\underline{(x-4)}}{3\underline{(x-4)}}\] any common factor?

OpenStudy (anonymous):

x-4

OpenStudy (ash2326):

yes, you're right. Did you understand?

OpenStudy (anonymous):

yes thank you